I need to convert 180 lbs of nitrogen/acre or 200 kg of nitrogen/hectare to the equivalent fertilizer application in an 8 liter container. Using a fertilizer with NPK 8-0-0.
Answer #1

Step 1. To calculate the required concentration of nitrogen in soil in Kg/L (from Kg/hectare)

Assumption : depth of soil = 20 cm = 0.2 m

We have 200 Kg Nitrogen/hectare ; 1 hectare = 10,000 m2. So, 200 Kg Nitrogen / 10, 000 m2;

if depth = 0.2 m, Volume = Area * height = 10, 000 * 0.2 = 2000 m3.

So, the required concentration is 200 Kg / 2000 m3= 0.1 Kg/m3.

We know that 1 m3 = 1000 L ; so the concentration in Kg/L = 0.1Kg/1000L = 10-4 Kg Nitrogen/L.

Step 2: Volume given = 8 L

so amount of nitrogen required = Concentration * volume = 10-4 * 8 = 8 * 10-4 Kg Nitrogen.

Step 3: Percentage of Nitrogen in the fertilizer bag = 8 %

So, [amount of fertilizer from the bag] * (8/100) = 8 * 10-4 Kg

[amount of fertilizer from the bag] = 8 * 10-4 Kg *(100/8) = 10-2 Kg = 10 g.